OpenStudy (anonymous):

Write the equation of the line that passes through (–4, 3) and is perpendicular to the line y=7/5x-3/5 A. 5x + 7y = 1 B. 7x + 5y = –13 C. 7x + 5y = –1 D. 3x + 5y = 7

OpenStudy (anonymous):

We can see that the equation y=7/5x-3/5 has a slope m=7/5 and a y-intercept b=-3/5. Now to find the slope of the perpendicular line, simply flip the slope m=7/5 to get m=5/7. Now change the sign to get m=-5/7. So the perpendicular slope is m=-5/7. Now let's use the point slope formula to find the equation of the perpendicular line by plugging in the slope and the coordinates of the given point (-4,3).

OpenStudy (anonymous):

note